# On solutions of reduced biquaternion equality constrained least squares problem and their relative forward error bound 22footnotemark: 2

### v1

@@ 3 An algebraic technique for complex solution of RBLSE problem
$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\;\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\;\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (25)
@@
$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (26)
@@
With the perturbed problem (25) and the bounds in (26), and using Theorem 3.1, the sensitivity analysis of the complex solution to the RBLSE problem (11) reduces to evaluating the sensitivity of the solution to the complex LSE problem (12). Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
4 An algebraic technique for real solution of RBLSE problem
$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\;\;\;\;B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$ (27)
$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\;\;\;\;\;D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ (28)
We will limit our discussion to the scenario where $m\geq n+d$, and both matrices $C$ and $C^{R}_{c}$ have full row rank. The RBLSE problem can then be stated as follows:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert AX-B\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;CX=D.$ (29)
Consider a real LSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}\;\;\;\textrm{subject to}\;\;\;C^{R}_{c}X=D^{R}_{c}.$ (30)
Let the QR factorization of $\left(C^{R}_{c}\right)^{T}$ be as follows:
where $\bar{Q}\in{\mathbb{R}}^{n\times n}$ is an orthonormal matrix and $\bar{R}\in{\mathbb{R}}^{4p\times 4p}$ is a nonsingular upper triangular matrix. Now, partition $A^{R}_{c}\bar{Q}$ as:
$A^{R}_{c}\bar{Q}=\left[\bar{P}_{1},\bar{P}_{2}\right],$ (32)
Theorem 4.1.
Consider the RBLSE problem outlined in (29) and the real LSE problem in (30), with notations specified in (31) and (32). For a matrix $X\in{\mathbb{R}}^{n\times d}$, $X$ is a real solution of the RBLSE problem (29) if and only if $X$ solves the real LSE problem (30). In this scenario, the unique solution with minimum norm $X$ can be expressed as:
Proof.
$\left\lVert A^{R}_{c}X-B^{R}_{c}\right\rVert_{F}=\textrm{min},\;\;C^{R}_{c}X=D^{R}_{c}.$ (34)
Using equations (2), (3), (5), along with Lemmas 2.1 and 2.3, we obtain
$=$ $\frac{1}{2}\left\lVert A^{R}X^{R}-B^{R}\right\rVert_{F}$
$=$ $\frac{1}{2}\left\lVert\left(AX-B\right)^{R}\right\rVert_{F}$
$=$ $\left\lVert AX-B\right\rVert_{F}.$
From (34), we obtain
and
$C^{R}X^{R}$ $=$ $D^{R},$
$(CX)^{R}$ $=$ $D^{R},$
$CX$ $=$ $D.$ (36)
By combining (35) and (36), we conclude that $X\in{\mathbb{R}}^{n\times d}$ is a real solution to the RBLSE problem (29), and vice versa. The expression for $X$ can be obtained by following the proof method of Theorem 3.1. ∎
$\widehat{A}=A+\Delta A,\;\widehat{B}=B+\Delta B,\;\widehat{C}=C+\Delta C,\;\textrm{and}\;\widehat{D}=D+\Delta D,$ (37)
where $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ represent the perturbations of the input data $A$, $B$, $C$, and $D$, respectively. We assume that the perturbation $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ are small enough to ensure that the perturbed matrices $\widehat{C}$ and $\widehat{C}^{R}_{c}$ retain full row rank. These perturbations are measured normwise by the smallest $\epsilon$ for which
$\left\lVert\Delta A\right\rVert_{F}\leq\epsilon\left\lVert A\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\left\lVert\Delta B\right\rVert_{F}\leq\epsilon\left\lVert B\right\rVert_{F},\;\textrm{and}\;\left\lVert\Delta D\right\rVert_{F}\leq\epsilon\left\lVert D\right\rVert_{F}.$ (38)
Let $\widehat{X}_{RL}$ be the real solution to the perturbed RBLSE problem
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\widehat{A}X-\widehat{B}\right\rVert_{F}\;\;\textrm{subject to}\;\;\widehat{C}X=\widehat{D},$ (39)
and let $\Delta X_{RL}=\widehat{X}_{RL}-X_{RL}$.
Theorem 4.2.
where
Proof.
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(\widehat{A}\right)^{R}_{c}X-\left(\widehat{B}\right)^{R}_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(\widehat{C}\right)^{R}_{c}X=\left(\widehat{D}\right)^{R}_{c}.$ (41)
$\left(\widehat{A}\right)^{R}_{c}=A^{R}_{c}+\left(\Delta A\right)^{R}_{c},\;\left(\widehat{B}\right)^{R}_{c}=B^{R}_{c}+\left(\Delta B\right)^{R}_{c},\;\left(\widehat{C}\right)^{R}_{c}=C^{R}_{c}+\left(\Delta C\right)^{R}_{c},\;\textrm{and}\;\left(\widehat{D}\right)^{R}_{c}=D^{R}_{c}+\left(\Delta D\right)^{R}_{c}.$
Thus, the perturbed real LSE problem (41) can be rewritten as:
$\min_{X\in{\mathbb{R}}^{n\times d}}\left\lVert\left(A^{R}_{c}+\left(\Delta A\right)^{R}_{c}\right)X-\left(B^{R}_{c}+\left(\Delta B\right)^{R}_{c}\right)\right\rVert_{F}\;\;\textrm{subject to}\;\;\left(C^{R}_{c}+\left(\Delta C\right)^{R}_{c}\right)X=\left(D^{R}_{c}+\left(\Delta D\right)^{R}_{c}\right).$ (42)
$\left\lVert\left(\Delta A\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta B\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{R}_{c}\right\rVert_{F},\;\left\lVert\left(\Delta C\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{R}_{c}\right\rVert_{F},\;\textrm{and}\;\left\lVert\left(\Delta D\right)^{R}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{R}_{c}\right\rVert_{F}.$ (43)
With the perturbed problem (42) and the bounds in (43), and using Theorem 4.1, the sensitivity analysis of the real solution to the RBLSE problem (29) reduces to evaluating the sensitivity of the solution to the real LSE problem (30). Consequently, the upper bound $U_{RL}$ for the relative forward error of the real solution to the RBLSE problem can be obtained from [1, Equation $4.11$]. ∎
@@ 5 Numerical Verification
Building on the previous discussions, this section presents numerical algorithms designed to find special solutions to the RBLSE problem. We also include numerical examples to validate these algorithms. Additionally, we assess the upper bound for the relative forward error of both complex and real solutions to the RBLSE problem.
@@
Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{C}_{c}$ have full row rank.
Output: $X$.
@@
Algorithm 2 For the real solution of the RBLSE problem (29)
Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that both matrices $C$ and $C^{R}_{c}$ have full row rank.
Output: $X$.
  1. 1.
  2. 2.
    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (32).
  3. 3.
    The real solution $X$ to the RBLSE problem (29) is given by (33).
Next, we provide examples to evaluate the effectiveness of the proposed algorithms. All computations were performed using a computer equipped with an Intel Core $i7-9700$ processor at $3.00$ GHz and $16$ GB of RAM, running MATLAB $R2021b$ software.
@@
Let $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}$. Let
$M_{i}$ $=$ $rand(m,n)+rand(m,n)\textit{{i}}\in{\mathbb{C}}^{m\times n},\;\textrm{for}\;i=1,2,$
$N_{i}$ $=$ $rand(m,d)+rand(m,d)\textit{{i}}\in{\mathbb{C}}^{m\times d},\;\textrm{for}\;i=1,2,$
$R_{i}$ $=$ $rand(p,n)+rand(p,n)\textit{{i}}\in{\mathbb{C}}^{p\times n},\;\textrm{for}\;i=1,2,$
$S_{i}$ $=$ $rand(p,d)+rand(p,d)\textit{{i}}\in{\mathbb{C}}^{p\times d},\;\textrm{for}\;i=1,2.$
Take $m=40t$, $n=6t$, $p=2t$, and $d=3$. Here, $t$ is an arbitrary number. We apply Algorithm 1 to determine the complex solution of the RBLSE problem. Let $X_{CL}$ be the complex solution of the RBLSE problem
$\min_{X,R_{c}}\left\lVert R_{c}\right\rVert_{F}\;\;\textrm{subject to}\;\;AX=B+R_{c},\;CX=D.$
Let $\epsilon_{1}=\left\lVert AX_{CL}-(B+R_{c})\right\rVert_{F}$ and $\epsilon_{2}=\left\lVert CX_{CL}-D\right\rVert_{F}$. In Table 1, we compute $\epsilon_{1}$ and $\epsilon_{2}$ for different values of $t$.
$t$ $\epsilon_{1}$ $\epsilon_{2}$
$1$ $8.5131\times 10^{-16}$ $2.0907\times 10^{-15}$
$3$ $7.0956\times 10^{-16}$ $2.5624\times 10^{-15}$
$5$ $1.0499\times 10^{-15}$ $3.7683\times 10^{-15}$
$7$ $1.2804\times 10^{-15}$ $4.5681\times 10^{-15}$
$9$ $1.4232\times 10^{-15}$ $8.1546\times 10^{-15}$
Table 1: Computational accuracy of Algorithm 1 for computing the complex solution of the RBLSE problem (11)
Table 1 shows that the errors $\epsilon_{1}$ and $\epsilon_{2}$ across different values of $t$ are consistently below $10^{-15}$. This indicates that Algorithm 1 is highly effective in determining the complex solution for the RBLSE problem (11).
@@
In Example 5.1, we introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively, to examine how these small perturbations affect the complex solution $X_{CL}$ of the RBLSE problem (11). These perturbations are measured normwise by the smallest $\epsilon$, as in (21). We define the exact relative forward errors for these solutions as $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$. Table 2 presents the calculated exact relative forward errors and the corresponding upper bound $U_{CL}$ (calculated using Equation 23) for these solutions across different values of $t$ and $\epsilon$.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
$1$ $7.0103\times 10^{-13}$ $1.3055\times 10^{-12}$ $2.7846\times 10^{-11}$
$7.0580\times 10^{-10}$ $1.3349\times 10^{-9}$ $2.8036\times 10^{-8}$
$8.9885\times 10^{-7}$ $1.81513\times 10^{-6}$ $3.5704\times 10^{-6}$
$5$ $7.1245\times 10^{-11}$ $1.1503\times 10^{-10}$ $9.5002\times 10^{-9}$
$8.9399\times 10^{-9}$ $2.0884\times 10^{-8}$ $1.1921\times 10^{-6}$
$1.0173\times 10^{-8}$ $1.5357\times 10^{-8}$ $1.3565\times 10^{-6}$
$9$ $7.1497\times 10^{-12}$ $1.2430\times 10^{-11}$ $1.3370\times 10^{-9}$
$6.6856\times 10^{-10}$ $1.2308\times 10^{-9}$ $1.2502\times 10^{-7}$
$1.0212\times 10^{-7}$ $2.5587\times 10^{-7}$ $1.9096\times 10^{-5}$
Table 2: Comparison of relative forward errors and their upper bounds for the complex solution of a perturbed RBLSE problem (11)
Table 2 shows that the exact relative forward errors of the complex solution to the RBLSE problem (11) are consistently lower than their respective upper bounds across various values of $t$ and $\epsilon$. This verifies the reliability of the derived upper bound $U_{CL}$ for the relative forward error.

### v2

@@ 4 An algebraic technique for complex solution of RBLSE problem
$\left(\widehat{A}\right)^{C}_{c}=A^{C}_{c}+\left(\Delta A\right)^{C}_{c},\quad\left(\widehat{B}\right)^{C}_{c}=B^{C}_{c}+\left(\Delta B\right)^{C}_{c},\quad\left(\widehat{C}\right)^{C}_{c}=C^{C}_{c}+\left(\Delta C\right)^{C}_{c},\quad\left(\widehat{D}\right)^{C}_{c}=D^{C}_{c}+\left(\Delta D\right)^{C}_{c}.$
@@
$\min_{X\in{\mathbb{C}}^{n\times d}}\left\lVert\left(A^{C}_{c}+\left(\Delta A\right)^{C}_{c}\right)X-\left(B^{C}_{c}+\left(\Delta B\right)^{C}_{c}\right)\right\rVert_{F}\quad\textrm{subject to}\quad\left(C^{C}_{c}+\left(\Delta C\right)^{C}_{c}\right)X=\left(D^{C}_{c}+\left(\Delta D\right)^{C}_{c}\right).$ (46)
@@
$\left\lVert\left(\Delta A\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert A^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta B\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert B^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta C\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert C^{C}_{c}\right\rVert_{F},\quad\left\lVert\left(\Delta D\right)^{C}_{c}\right\rVert_{F}\leq\epsilon\left\lVert D^{C}_{c}\right\rVert_{F}.$ (47)
@@
With the perturbed problem (46) and the bounds in (47), and using Theorem 4.1, the sensitivity analysis of the complex solution to the RBLSE problem (33) reduces to evaluating the sensitivity of the solution to the complex LSE problem (34). Consequently, the upper bound $U_{CL}$ for the relative forward error of the complex solution to the RBLSE problem can be obtained from [2, Equation $4.11$]. ∎
@@ 5 Numerical Verification
In this section, we present numerical algorithms for computing two special solutions (real and complex) of the RBLSE problem. Each algorithm is complemented with a detailed computational complexity analysis through a step-by-step flop count. Additionally, we validate these algorithms using numerical examples to assess their performance and reliability.
Algorithm 1 For the real solution of the RBLSE problem (16)
Input: $A=A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that matrix $C^{R}_{c}$ has full row rank.
Output: $X\in{\mathbb{R}}^{n\times d}$.
  1. 1.
  2. 2.
    Partition the matrix $A^{R}_{c}\bar{Q}$ as shown in (19).
  3. 3.
    The real solution $X$ to the RBLSE problem (16) is given by (20).
@@
Input: $A=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n}$, $B=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d}$, $C=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n}$, and $D=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$ We assume $m\geq n+d$ and that matrix $C^{C}_{c}$ has full row rank.
Output: $X\in{\mathbb{C}}^{n\times d}$.
@@
Next, to assess the performance of the algorithms, a step-by-step flop count is presented. This detailed analysis highlights the computational cost at each stage, providing insights into the efficiency of the algorithm. First, to analyze the efficiency of the real solution algorithm, a detailed step-by-step flop count is provided below.
Step Description Flop Count
1 QR decomposition of $\left(C^{R}_{c}\right)^{T}$ $O(32np^{2})$
2 Compute $A^{R}_{c}\bar{Q}$ $O(8mn^{2}-4mn)$
3 Partition $A^{R}_{c}\bar{Q}$ into $\bar{P}_{1}$ and $\bar{P}_{2}$ $O(1)$
4 Compute $\bar{P}_{2}^{\dagger}$ $O\big{(}24m(n-4p)^{2}+10(n-4p)^{3}\big{)}$
5 Solve $\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(16p^{2}d)$
6 Compute $\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(32mpd-4md)$
7 Compute $B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c}$ $O(4md)$
8 Compute $\bar{P}_{2}^{\dagger}(B^{R}_{c}-\bar{P}_{1}\left(\bar{R}^{T}\right)^{-1}D^{R}_{c})$ $O(8m(n-4p)d)$
9 Compute $X$ $O(2n^{2}d)$
Total Flop Count $O\big{(}32np^{2}+8mn^{2}-4mn+24m(n-4p)^{2}+10(n-4p)^{3}+16p^{2}d+32mpd-4md+4md+8m(n-4p)d+2n^{2}d\big{)}$
Table 1: Flop count for the computational steps to find the real solution of the RBLSE problem.
To analyze the efficiency of the complex solution algorithm, a detailed step-by-step flop count is provided below. This analysis emphasizes the additional computations required due to the complex structure.
Step Description Flop Count
1 QR decomposition of $\left(C^{C}_{c}\right)^{H}$ $O(32np^{2})$
2 Compute $A^{C}_{c}\widetilde{Q}$ $O(16mn^{2}-4mn)$
3 Partition $A^{C}_{c}\widetilde{Q}$ into $\widetilde{P}_{1}$ and $\widetilde{P}_{2}$ $O(1)$
4 Compute $\widetilde{P}_{2}^{\dagger}$ $O\big{(}48m(n-2p)^{2}+40(n-2p)^{3}\big{)}$
5 Solve $\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(16p^{2}d+10pd)$
6 Compute $\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(32mpd-4md)$
7 Compute $B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c}$ $O(4md)$
8 Compute $\widetilde{P}_{2}^{\dagger}(B^{C}_{c}-\widetilde{P}_{1}\left(\widetilde{R}^{H}\right)^{-1}D^{C}_{c})$ $O(16m(n-2p)d)$
9 Compute $X$ $O(8n^{2}d)$
Total Flop Count $O\big{(}32np^{2}+16mn^{2}-4mn+48m(n-2p)^{2}+40(n-2p)^{3}+16p^{2}d+10pd+32mpd-4md+4md+16m(n-2p)d+8n^{2}d\big{)}$
Table 2: Flop count for the computational steps to find the complex solution of the RBLSE problem.
Next, we provide examples to evaluate the effectiveness of the proposed algorithms. All computations were performed using a computer equipped with an Intel Core $i9-12900K$ processor at $3200$ MHz and $32$ GB of RAM, running MATLAB $R2024b$ software.
The following example validates the efficiency of the algorithms by comparing the real and complex solutions in terms of computational accuracy and CPU runtime.
@@ 9 Compute $X$ $O(8n^{2}d)$
To evaluate the accuracy and performance of the proposed algorithms for solving the RBLSE problem, we use random data matrices of varying sizes. The problem matrices are defined as follows:
$A$ $=$ $A_{0}+A_{1}\textit{{i}}+A_{2}\textit{{j}}+A_{3}\textit{{k}}=M_{1}+M_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times n},\quad B=B_{0}+B_{1}\textit{{i}}+B_{2}\textit{{j}}+B_{3}\textit{{k}}=N_{1}+N_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{m\times d},$
$C$ $=$ $C_{0}+C_{1}\textit{{i}}+C_{2}\textit{{j}}+C_{3}\textit{{k}}=R_{1}+R_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times n},\quad D=D_{0}+D_{1}\textit{{i}}+D_{2}\textit{{j}}+D_{3}\textit{{k}}=S_{1}+S_{2}\textit{{j}}\in\mathbb{Q}_{\mathbb{R}}^{p\times d}.$
The real matrix components $A_{i},B_{i},C_{i},$ and $D_{i}$ are generated using random values:
$A_{i}$ $=rand(m,n)\in{\mathbb{R}}^{m\times n},\quad i=0,1,2,3,$
$B_{i}$ $=rand(m,d)\in{\mathbb{R}}^{m\times d},\quad i=0,1,2,3,$
$C_{i}$ $=rand(p,n)\in{\mathbb{R}}^{p\times n},\quad i=0,1,2,3,$
$D_{i}$ $=rand(p,d)\in{\mathbb{R}}^{p\times d},\quad i=0,1,2,3.$
The matrix dimensions are determined by $m=30t$, $n=10t$, $p=2t$, and $d=2$, where $t$ is a positive integer. We apply Algorithms 1 and 2 to compute the real and complex solutions of the RBLSE problem, respectively.
  1. (a)
    Let $X_{RL}$ be the real solution of the RBLSE problem:
    $\min_{X,R_{r}}\|R_{r}\|_{F}\quad\text{subject to}\quad AX=B+R_{r},\quad CX=D.$
    The accuracy of $X_{RL}$ is measured using the following errors:
    $\epsilon_{1}=\log_{10}(\|AX_{RL}-(B+R_{r})\|_{F}),\quad\epsilon_{2}=\log_{10}(\|CX_{RL}-D\|_{F}).$
  2. (b)
    Let $X_{CL}$ be the complex solution of the RBLSE problem:
    $\min_{X,R_{c}}\|R_{c}\|_{F}\quad\text{subject to}\quad AX=B+R_{c},\quad CX=D.$
    The accuracy of $X_{CL}$ is measured using:
    $\epsilon_{3}=\log_{10}(\|AX_{CL}-(B+R_{c})\|_{F}),\quad\epsilon_{4}=\log_{10}(\|CX_{CL}-D\|_{F}).$
To thoroughly evaluate the computational performance, the CPU times for each solution are recorded and averaged over $50$ trials to minimize variability. The average CPU times (in seconds) for computing $X_{RL}$ and $X_{CL}$ are denoted by $t_{r}$ and $t_{c}$, respectively.
$t$ $\epsilon_{1}$ $\epsilon_{2}$ $\epsilon_{3}$ $\epsilon_{4}$
$1$ $-15.1057$ $-15.0513$ $-15.5746$ $-15.2797$
$3$ $-15.1474$ $-14.5131$ $-15.4137$ $-14.7818$
$5$ $-14.9570$ $-14.4140$ $-15.3090$ $-14.4512$
$7$ $-14.8277$ $-14.2473$ $-15.3418$ $-14.5228$
$9$ $-14.8800$ $-14.0209$ $-15.3164$ $-14.2537$
Table 3: Accuracy of Algorithms 1 and 2 in computing the real and complex solutions of the RBLSE problem.
Table 3 demonstrates that the two algorithms achieve high accuracy, with the computational errors consistently below $-14$ for all tested values of $t$. This highlights the accuracy of the proposed methods.
Figure 1: Comparison of CPU time for computing real and complex solution of RBLSE problem
Figure 1 compares the average CPU times for the two algorithms. The results show that the algorithm for computing real solution takes less time compared to the algorithm for computing the complex solution of the RBLSE problem. This is consistent with the theoretical complexity of the algorithms.
The following example compares the theoretical upper bounds with the actual relative forward errors for the real and complex solutions of the RBLSE problem, demonstrating the reliability of the proposed bounds.
@@
Building on Example 5.1, we now introduce random perturbations $\Delta A$, $\Delta B$, $\Delta C$, and $\Delta D$ to the input matrices $A$, $B$, $C$, and $D$, respectively. These perturbations are used to examine the sensitivity of the RBLSE problem to small changes in the input data. Specifically, we evaluate how these perturbations affect the computed real solution $X_{RL}$ and the complex solution $X_{CL}$. The magnitude of these perturbations is measured normwise by the smallest $\epsilon$, which ensures the changes are controlled and small.
To quantify the impact of the perturbations, we compute the exact relative forward errors for each solution as follows:
$\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}\quad\text{and}\quad\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}.$
Here, $\widehat{X}_{RL}$ and $\widehat{X}_{CL}$ denote the computed solutions of the perturbed RBLSE problem, while $X_{RL}$ and $X_{CL}$ are the original solutions.
In addition to the exact relative forward errors, we compute the corresponding upper bounds $U_{RL}$ and $U_{CL}$ for these errors. These bounds are derived using equations (27) and (44), respectively, which provide theoretical estimates of the maximum possible errors.
Table 4 presents the exact relative forward errors and their corresponding upper bounds for various values of $t$ and $\epsilon$, allowing us to evaluate the relationship between the perturbation magnitude and the stability of the solutions.
$t$ $\epsilon$ $\frac{\|\widehat{X}_{RL}-X_{RL}\|_{F}}{\|X_{RL}\|_{F}}$ $U_{RL}$ $\frac{\|\widehat{X}_{CL}-X_{CL}\|_{F}}{\|X_{CL}\|_{F}}$ $U_{CL}$
$1$ $1.2957\times 10^{-13}$ $4.5138\times 10^{-13}$ $7.6470\times 10^{-12}$ $3.8724\times 10^{-13}$ $5.6680\times 10^{-12}$
$1.3279\times 10^{-10}$ $4.3088\times 10^{-10}$ $7.8370\times 10^{-9}$ $1.8961\times 10^{-10}$ $5.8088\times 10^{-9}$
$1.2770\times 10^{-7}$ $3.0154\times 10^{-7}$ $7.5367\times 10^{-6}$ $2.8896\times 10^{-7}$ $5.5863\times 10^{-6}$
$5$ $1.2312\times 10^{-11}$ $6.1760\times 10^{-11}$ $4.7674\times 10^{-9}$ $3.1584\times 10^{-11}$ $1.7960\times 10^{-9}$
$1.5256\times 10^{-9}$ $9.3350\times 10^{-9}$ $5.9071\times 10^{-7}$ $4.8613\times 10^{-9}$ $2.2254\times 10^{-7}$
$1.4801\times 10^{-8}$ $7.9283\times 10^{-8}$ $5.7313\times 10^{-6}$ $3.3857\times 10^{-8}$ $2.1592\times 10^{-6}$
$9$ $1.2645\times 10^{-12}$ $6.5592\times 10^{-12}$ $4.5359\times 10^{-10}$ $4.0664\times 10^{-12}$ $2.4556\times 10^{-10}$
$1.2593\times 10^{-10}$ $6.6261\times 10^{-10}$ $4.5171\times 10^{-8}$ $4.0808\times 10^{-10}$ $2.4454\times 10^{-8}$
$1.5340\times 10^{-7}$ $7.1997\times 10^{-7}$ $5.5023\times 10^{-5}$ $3.8176\times 10^{-7}$ $2.9788\times 10^{-5}$
Table 4: Comparison of relative forward errors and their upper bounds for the real and complex solutions of a perturbed RBLSE problem.
Table 4 illustrates that the exact relative forward errors for the real and complex solutions of the RBLSE problem are consistently smaller than their respective upper bounds $U_{RL}$ and $U_{CL}$, for all considered values of $t$ and $\epsilon$. These results validate the accuracy and reliability of the derived upper bounds.
The following example evaluates the accuracy of the proposed algorithms for computing the real and complex solutions of the RBLSE problem.